OpenStudy (anonymous):

Which equation would best help solve the following problem? Holly has a rectangular garden that measures 12 m wide by 14 m long. She wants to increase the area to 255 m2 by increasing the width and length by the same amount. What will be the dimensions of the new garden?

OpenStudy (anonymous):

A.255 = (12)(14) B.255 = (12 + x2)(14) C.255 = (12 + x)(14 + x) D.255 = (12 – x)(14 – x)

OpenStudy (anonymous):

@SolomonZelman @iambatman

OpenStudy (solomonzelman):

What do you think ?

OpenStudy (solomonzelman):

If you are increasing each dimension by x, then ...

OpenStudy (anonymous):

Well, B? It is increasing by 2x

OpenStudy (acxbox22):

not 2x, x^2 the new are is 255 your increasing 12 by x and 14 by x when multiplied they equal the new area

OpenStudy (solomonzelman):

EACH dimension increases by x.

OpenStudy (anonymous):

Oh so * each by +x?

OpenStudy (anonymous):

Holly has a rectangular garden that measures 12 m wide by 14 m long. She wants to increase the area to 255 m2 by increasing the width and length by the same amount. What will be the width (shorter dimension) of the new garden? A.14 m wide B.17 m wide C.16 m wide D.15 m wide

OpenStudy (anonymous):

Going with C for the first one

OpenStudy (acxbox22):

good job, C is right for number 1

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

(12 + x)(14 + x) = 255 x^2 + 26x + 168 = 255 x^2 + 26x - 87 = 0 (x + 29)(x - 3) x = -29, 3 x = 3 (increased by) so new dimensions is 15, 17 length is 17 m

OpenStudy (acxbox22):

to do number 2 just solve the quadratic equation 255=(x+14)(x+12) 255=x^2+26x+168 0=x^2+26-87 0=(x+29)(x-3) 0=x+29 x=-29 and x=3 x=3 is the only sensible solution 3+12=15 so 15 m wide

OpenStudy (anonymous):

Ty

OpenStudy (anonymous):

Is 17 the Longer length?

OpenStudy (anonymous):

@acxbox22

OpenStudy (acxbox22):

yes

OpenStudy (anonymous):

Thanks